OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations?

OpenStudy (anonymous):

4x + 2y = 6 x – y = 3

OpenStudy (anonymous):

−1 5 −2 2

OpenStudy (anonymous):

Solve by using elimination.

OpenStudy (anonymous):

4x + 2y = 6 x – y = 3 double the second one get \[4x+2y=6\] \[2x-2y=6\]

OpenStudy (anonymous):

and them up in a column, get \[6x=12\] and so \(x=2\)

OpenStudy (anonymous):

I thought you add the two equations together?